A service station wishes to mix 40 L of a 78% antifreeze solution. How many litres of a 75% solution and a 90% solution should be used in forming the mixture? ( using two unknowns, write two equations and solve.)
Answer by checkley71(8403)

.9X+.75(40-X)=40*.78
.9X+30-.75X=31.2
.15X=31.2-30
.15X=1.2
X=1.2/.15
X=8L OF 90% IS NEEDED.
40-8=32L OF 75% IS NEEDED.
PROOF
.9*8+.75*32=40*.78
7.2+24=31.2
31.2=31.2